A truck covers 40.0 m in 8.50 s while smoothly slowing down to a final speed of 2.80 m/s. (a) Find its original speed. (b) Find its acceleration.
1
Expert's answer
2020-10-07T10:17:55-0400

B) Since "V_1=V_0-a\times t" , where "V_1" is the final velocity, "V_0" is the initial velocity, then "V_0=V_1+a\times t" ,

and since "S =\frac {V_1^2-V_0^2}{2\times a}" ,

"S=\frac {V_1^2-V_1^2+2\times V_1\times a \times t + a^2 \times t^2}{2\times a}"

"a=\frac{2\times S- 2\times V_1 \times t }{t^2}=\frac{2\times 40- 2\times 2.8 \times 8.5}{8.5^2}\approx0.45" m/s"^2"

а) "V_0=V_1+a\times t=2.8+0.45\times 8.5=6.625" m/s